okay. my question is 4m+7=43.8
4m=43.8-7
4m=36.8 NOW what do i do, To finish the answer
because i don't know what to do with the 4m
HELP !!!!!!!!!!!

Answer by nerdybill(7384)

When you get here:
4m=43.8-7
4m=36.8
to isolate 'm' we divide both sides by 4:
(4m)/4 = (36.8/4)
notice the 4s cancel on the left:
m = 36.8/4
m = 9.2